a bike rental shop charges a one-time charge of $8 plus an hourly fee to rent a bike. Dan paid $24.50 to rent a bike for 5 1/2 hours.Find the bike shop's hourly fee dollars.

i realy need help!!!!!!!!!!!

Answer by nerdybill(5404)

a bike rental shop charges a one-time charge of $8 plus an hourly fee to rent a bike. Dan paid $24.50 to rent a bike for 5 1/2 hours.Find the bike shop's hourly fee dollars.
.
Let x = hourly rate
then from "Dan paid $24.50 to rent a bike for 5 1/2 hours." we get:
24.50 = 8 + 5.5x
.
Now, we solve for x:
24.50 = 8 + 5.5x
16.50 = 5.5x
16.50/5.5 = x
3 dollars/hour = x
.
Hourly fee: $3 dollars